OpenStudy (anonymous):

Bob and James are finishing the roof of a house. Working alone, Bob can shingle the roof in 14 hours. James can shingle the same roof in 18 hours. How long will it take them working together to shingle the roof? Round your answer to the nearest hundredth if necessary.

OpenStudy (raden):

working together (time) = (t1 x t2)/(t1 + t2) = (14 x 18)/(14 + 18) = 252/32 = 7.875 h with t1 = bob's time and t2 = James's time